OpenStudy (anonymous):

Find the area of the following rectangle. a = 12 ft. b = 14 ft. a = ____sq. ft.

OpenStudy (anonymous):

168

OpenStudy (ajprincess):

Area=length*width =b*a =12*14 =168

OpenStudy (mayankdevnani):

@Claire4christ tag

OpenStudy (mayankdevnani):

very fast calculation

OpenStudy (anonymous):

168 sq ft right we don't have to convert them ??

OpenStudy (mayankdevnani):

no @Claire4christ

OpenStudy (ajprincess):

u hav to convert if u r asked to give the answer in cm^2 or m^2

OpenStudy (anonymous):

ok